if f(x)=x^2+3 and g(x)=3x-1
then find the following

1d. (fg)(x)

1e. fg(-1)

1f. (f/g)(x)
1
Expert's answer
2012-12-20T09:06:06-0500

1d. (fg)(x)=(x^2+3)(3x-1)=3x^3-x^2+9x-3
1e. (fg)(-1)=3(-1)^3-(-1)^2+9(-1)-3=-3-1-9-3=-16
1f. (f/g)(x)=(x^2+3)/(3x-1)